OpenStudy (anonymous):

A truck undergoes a constant acceleration of 2.0 meters/second2 on a highway that slopes upward at an angle of 30.0° to the horizontal. What is the magnitude of the horizontal component of the acceleration?

OpenStudy (souvik):

a*cos(theta)

OpenStudy (anonymous):

i dont get this fromeula.. as you can see im a loser at science its my worst subject..

OpenStudy (souvik):

do you know what is cos(theta)=?

OpenStudy (anonymous):

nope/:

OpenStudy (souvik):

okay...let me explain..

OpenStudy (souvik):



OpenStudy (souvik):

cos(theta )=AB/AC and sin(theta)=CB/AC

OpenStudy (souvik):

now AC=a then cos(theta)=AB/a or AB=a*cos(theta)

OpenStudy (souvik):

in this question a and theta is known to you...need to know AB

OpenStudy (anonymous):

now i get it (:

OpenStudy (souvik):

:)